OpenStudy (anonymous):

Help pleasee! When a golfer plays any hole, he will take 3,4,5,6, or 7 strokes with probabilities of 1/10, 1/5, 2/5, 1/5, and 1/10, respectively. He never takes more than 7 strokes. Find the probability of: A) Scoring 4 on each of the first three holes. B) Scoring 3,4 and 5 (in that order( on the first three holes. C) Scoring a total of 28 for the first four holes. D) Scoring a total of 10 for the first three holes. Honestly, any help is greatly appreciated; my teacher's been trying to explain over and over and I still get confused :(

OpenStudy (kropot72):

(A) The probability of scoring 4 on each of the first three holes is \[(\frac{1}{5})^{3}\] (B) The probability of scoring 3, 4 and 5 on the first three holes is \[\frac{1}{10}\times \frac{1}{5}\times \frac{2}{5}\] There are 3! permutations of the three scores of which only one is required. Therefore the probability of scoring 3,4 and 5 (in that order) on the first three holes is \[\frac{1}{6}\times \frac{1}{10}\times \frac{1}{5}\times\frac{2}{5}\]